# Predicting potentially unfair clauses in Chilean terms of services with natural language processing

### v1

@@ Abstract
This study addresses the growing concern of information asymmetry in consumer contracts, exacerbated by the proliferation of online services with complex Terms of Service that are rarely even read. Even though research on automatic analysis methods is conducted, the problem is aggravated by the general focus on English-language Machine Learning approaches and on major jurisdictions, such as the European Union. We introduce a new methodology and a substantial dataset addressing this gap. We propose a novel annotation scheme with four categories and a total of 20 classes, and apply it on 50 online Terms of Service used in Chile. Our evaluation of transformer-based models highlights how factors like language- and/or domain-specific pre-training, few-shot sample size, and model architecture affect the detection and classification of potentially abusive clauses. Results show a large variability in performance for the different tasks and models, with the highest macro-F1 scores for the detection task ranging from 79% to 89% and micro-F1 scores up to 96%, while macro-F1 scores for the classification task range from 60% to 70% and micro-F1 scores from 64% to 80%. Notably, this is the first Spanish-language multi-label classification dataset for legal clauses, applying Chilean law and offering a comprehensive evaluation of Spanish-language models in the legal domain. Our work lays the ground for future research in method development for rarely considered legal analysis and potentially leads to practical applications to support consumers in Chile and Latin America as a whole.
@@
Consumer protection law, Unfair clauses, Natural language processing, Machine learning, Neural networks
@@ 1 Introduction
In the digital age, where a growing number of products and services are offered through online platforms, Terms of Service (ToS) have become increasingly relevant, since they usually constitute the step where the provider and the user become legally bound. However, ToS are long, difficult to read, sometimes not even available in the mother tongue of the user and, at the end of the day, pointless to review. Even if one could and would want to spent time reading them in order to use the service, there is no alternative but to accept them. It is not possible to change or negotiate a disadvantageous or unfair clause[19]. These problems of information asymmetry and rational apathy are present in most modern consumer contracts, but now they are amplified by the number and speed they occur through digital platforms. This constitutes a challenge for the regulator, the surveillance agencies and the consumers [40].
In this context, the use of Artificial Intelligence in the legal field could be of tremendous use, because it could automate the revision of these ToS and other consumer contract, alerting the user and national services of the existence of potentially unfair clauses [36]. Recent developments have already led to promising systems, however, with few exceptions [59], these have mainly focused on the European Union [40, 8, 21, 52, 6, 23, 28]. Methods developed for other legal systems and languages, such as South American legislation developed in Spanish, are also necessary and should be included in the global research community’s attention.
Hence, our research objectives focus on the development of Machine Learning (ML) for the automated identification of potentially abusive clauses in Chilean online ToS. Hereby, we build on work conducted on European online ToS [40, 52, 13, 23, 28] and update the methodology to recent developments in Natural Language Processing (NLP) [47, 22, 63, 41].
@@
  • We present a classification system categorizing 20 potentially abusive clauses into three groups: 6 illegal, 6 dark, and 8 gray.
  • We introduce the Chilean ToS dataset, comprising 50 legal documents in Spanish with a total of 6,523 annotated clauses. From these, we generate six datasets of the tasks detection or classification, resulting in a corpus of in total 15,201 clauses. These datasets are imbalanced and either binary or multi-class/-label classification.
@@
The rest of the paper is structured as follows. The next section formulates the problem from the Chilean legal perspective. Section 3 presents the proposed annotation system and introduces the Chilean Terms of Services corpus and its datasets. Section 4 explains ML methods and Section 5 describes the experimental setup and reports the results of the study that we conducted. Section 6 discusses relevant prior research and provides an outlook, and Section 7 concludes.
@@ 2 Problem Formulation
Given the asymmetry of information and the rational apathy issues prevalent in modern consumer contracts, there is a risk that a provider could include unfair clauses in them. Legal systems refer to these types of clauses in various ways. In Chile, the Consumer Protection Law Act (Law Number 19,496, "Ley de Protección al Consumidor", recently updated and modified by the "Decreto con Fuerza de Ley Numero 3" from 2021, from now on LPC) does not provide a definition of unfair clauses. Instead, it outlines the requirements that every consumer contract must fulfill and lists clauses that will be deemed invalid if introduced and that could potentially nullifying the entire contract (Article 16 of the LPC). The Chilean National Service for Consumer Protection ("Servicio Nacional del Consumidor", SERNAC) has defined abusive clauses as those imposed by the provider of the goods or services in violation of the principles of equality and good faith, that result in a significant imbalance of contractual duties, in detriment of the weakest part, the consumer [55]. Examples of these clauses are the ones that allow the unilateral termination of the contract at the mere discretion of the provider, that is, without expression of cause or without the fault of the consumer, that include disclaimers of liability even when the damage was cause by the supplier, or that cause a reversal of the burden of proof, among many others unfair clauses.
@@
The core of these types of clauses lies in the power imbalance between the parties involved in the contract. One way to restore that balance is to report the existence of unfair clauses and eventually pursue legal action against provider who use them. However, the first requirement for this is that consumers are aware of them. According to a survey conducted by SERNAC, over half of consumers are unable to identify potentially unfair clauses in contracts, while a third are completely unaware of their existence  [^1] https://www.sernac.cl/portal/604/w3-article-64918.html, (March 1st, 2022). This situation is exacerbated even further in case of ToS, that are designed to be accepted just by one click in matter of seconds [40].
@@
In this context, specialized ML methods could be used as a preventive tool by the general public to check if a ToS or other contracts include potentially unfair clauses. A task that would be arduous and time-consuming for a human could be swiftly accomplished by the model. This model could automatically review the entire extension of the terms and raise an alarm if if detects any clause that appears to be abusive. The primary objective of this research is to develop such methods.
@@
It is crucial to recognize that certain clauses can be interpreted in various ways, depending on the systematic analysis of the contract and the factual context in which they are applied. This causes interpretation problems that affect even specialized lawyers, thereby posing a substantial challenge for ML algorithms. Furthermore, the judicial authority is responsible for determining whether a clause is unfair or illegal. They will impose the appropriate consequences, which may include the annulment of the clause or contract, and eventually, the compensation of the suffered damages. In that regard, until the Supreme Court has settled the discussion, a clause is only potentially unfair. In other words, neither an AI method, nor a lawyer or nor even SERNAC can definitively determine whether a clause is unfair. Only the Supreme Court possesses the authority to affirm or deny with legal certainty the nature of a clause. Thus, our proposed ML methods and their results should be understood as a reference to guide further actions, but are not intended as a replacement of the professional and technical advise. However, they can still serve as a very useful preventive tool that can raise awareness among consumers about the existence of unfair clauses, and may support the work of surveillance agencies and lawyers who deal with such clauses.
3 Potentially Unfair Clauses
@@
In this study we analyzed 50 Spanish-language ToS from companies that provide services in Chile  [^2] Academia.edu, Airbnb, Amazon, App Copec, Apple, Badoo, Bluexpress, Booking, Box, Canva, Despegar, Dropbox, Ebay, Edreams, Evernote, Facebook, Fitbit, Google, Headspace, Instagram, Linkedln, Mercadolibre, Microsoft, MyHeritage, Nespresso, Netflix, Nike, Nintendo, Paris, Pokemon GO, Pullman Bus, Rappi, Ripley, Rovio, Skype, Skyscanner, Snapchat, Spotify, Starbucks, Tenpo, Tinder, TripAdvisor, Uber, Vimeo, WhatsApp, Wild Foods, World of Warcraft, X (Twitter), Yahoo, YouTube. In total, this corresponds to 6,523 clauses, of which $20.3\%$ were identified as potentially unfair under at least one category. That means, they were clauses that presented more than one reason that deemed them abusive, so the level of unfairness is actually underrepresented. That is a surprisingly high level of unfair clauses, since one could expect that standardized contracts used by internationally renown corporations should comply with the law. Again this result reaffirm the necessity of developing tools that alert consumers of the existence of potentially unfair clauses.
@@
From the about 20% of clauses that were deemed unfair by our legal expert annotators, some of them were quite common. For example we found 156 clauses in the corpus that limit the provider’s liability (ltd). On the other hand, there are extremely rare cases, such as the 10 clauses in the corpus that permit the provider the unilateral termination of the contract (ter). Furthermore, several clauses may be potentially abusive in multiple ways. For instance, out of the 117 clauses that impose external costs on the consumer (des risk), 34 were also against the good faith (bfe). In the following sections, we describe our proposed categorization system and its type of potentially abusive clauses, the annotated clauses and finally the corpus statistics.
@@ 3.1 Categorization System
For reviewing the ToS we developed our own categorization system based on three factors. First, we used the legal classification Chilean doctrine has made of Abusive Clauses [19, 37, 4, 43, 51, 55]. Second, we reviewed the classification used in previous related work [40, 52], and finally, we took the optimal methodology to identify the clauses into consideration. For this work, we define four categories of potentially unfair clauses for Chilean law: formal requirements, and illegal, dark or gray clauses. Of these four, only the last three are predicted in our experiments, as the next section will clarify.
@@ 3.2 Formal Requirements of Clauses
Some norms of the LPC, like Articles 17 and 17A, specify different formal requirements that contracts must fulfill regardless of their type or content. For example, the minimum size of the letter they should use, the language in which they should be written, the obligation of providing a print copy of the contract, among others. Under these requirements many of the ToS were at fault at very basic elements, for example, some of them were not available in Spanish  [^3] Those were excluded from this study that focus on this language, or were not printable.
These types of infringements of the law, even though relevant to consumers, are not of special interest in our current research, since their identification is fairly straightforward and no specialized model is required for it. Instead, more traditional text parsers could be used to analyze, for example, the language of the ToS, its HTML formatting and font size or the blocking of any print functionality via JavaScript. Therefore, they were excluded from the experimental analysis of this work, that focuses on the development and review of the performance of ML methods.
@@ 3.3 Illegal Clauses
Illegal Clauses refer to provisions that go against the explicit order of a norm in detriment of the consumer. Therefore, even though they are included in a contract or ToS, they are not effective, and should be considered as "not written", according to Articles 1,467 and 1,682 of the Civil Code ("Código Civil", CC) regarding the effect of annulment of clauses.
This category is not one used by the Chilean doctrine or previous related work, but is of our own creation. We decided to separate this type of clauses due to two reasons. First, because they constitute a relevant case of potentially unfair clauses (about 45% of them), and, secondly, because their characteristics make them both easy to recognize and at she same time uniquely challenging for an AI system. When the ML model was trained on the annotated dataset, it was able to successfully predict the lawfulness of a clause. However, a contractual provision could potentially violate any legal rule of the Chilean system, and therefore, it can be extremely difficult to identify and group the applicable norms. For example, during the annotation of ToS, we found that a quite specific law was infringe by a ToS, Law Number 20,720 regarding the bankruptcy regime and insolvency procedures. This was done due to the annotators’ expert knowledge, but could not have been otherwise expected in the review of the ToS. Like that maybe other clauses were unlawful, but were not detected by the human and thus the ML model was not trained with the required data.
In the process, we noticed that some illegalities were more common than others, so we proceed to subdivided them into their own classes. For example, the clauses that contravene provisions of the contractual liability regime of the CC, that included infringements in the rules for the determination of damages or the binding nature of contracts. Another usual illegal clauses were the ones that altered the jurisdiction of courts or prevented the consumer from acceding to one, but instead forced them to use informal conflict mechanism resolutions, violating procedural rules of the LPC, the Civil Procedure Code ("Código de Procedimiento Civil", CPC) and the Organic Code of Courts ("Código Orgánico de Tribunales, COT). Since many ToS come from international corporations, clauses referencing foreign law, that are not applicable in Chile (and sometimes even in direct conflict with national provisions), were also identified as illegal, in order to ensure local consumer rights are upheld.
@@ 3.4 Dark Clauses
Dark Clauses are the ones that, even though they are not directly illegal, they are manifestly unfair. They are caused by the great power imbalance the provider holds, since the consumer can only agree with them or not sign the contract, i.e., not use the service. The provisions included in letters a), b), c) d), e) and h) of Article 16 LPC refer to this type of clauses [4]. These are, for example, the unilateral termination of the contract at the mere discretion of the provider; the arbitrary modification of the elements of the contract, including an unilateral price increase; preemptive limitation of liability clauses; alteration of the burden of proof.
A commonality among these clauses is how evidently imbalance they are, and for that reason the Chilean legislator presumes their unfairness. This means that, even though their identification as abusive still needs to be done by a court, the burden of proof lies with the supplier and not on the consumer. In other words, the provider that incorporated a dark clause into their contract needs to prove that, despite their appearance, they are not actually unfair. This constitutes a important advantage at the moment of presenting a claim, but a requirement for this step is that consumers are aware of their existence and unfairness. According to the studies conducted by SERNAC (see footnote 1), this is not a simple task for most people. Perhaps it is because of this that, despite the legislator identifying them as unfair, they continue to be widely used by providers.
Dark Clauses present three special characteristics that might be beneficial for ML. First, the above mention letters of Article 16 LPC constitute a well define list of clauses, that stands in contrast with the countless sources of unlawfulness that can make a clause illegal. Closed-set classification problems, such as this, are generally considered easier to solve than open-set classification like the Illegal Clauses problem. For instance, there is no need for open-set recognition [60]. However, the challenge of the variability of examples for one type of potentially unfair clause remains, i.e., they can be reformulated in different ways that are not recognizable anymore, for instance, using synonyms or a different grammatical structure. Second, since they are presumed unfair, there is less error of interpretation and they are relatively clear. In contrast, Gray Clauses are more ambiguous, see Sec. 3.5. While the final decision is with the courts, their identification via ML would help raise awareness of the clauses existence among consumers. Third, since many Dark Clauses appear relatively frequently in the ToS that we reviewed, and thus, their patterns and specific formulations repeat as well, ML models may actually learn to predict these more easily. Conversely, the clauses that are less represented in the data may be harder to predict accurately.

### v2

# Predicting potentially abusive clauses in Chilean terms of services with natural language processing

@@ Abstract
This study addresses the growing concern about the inclusion of abusive clauses in consumer contracts, exacerbated by the proliferation of online services with complex Terms of Service that are rarely read. Even though research on automatic analysis methods is conducted, the difficulty of detecting such clauses is aggravated by the general focus on English-language Machine Learning approaches and on major jurisdictions, such as the European Union. We introduce a new methodology and a substantial Spanish-language dataset addressing this gap. We propose a novel annotation scheme with four categories and 20 classes and apply it to 50 online Terms of Service used in Chile. Our evaluation of transformer-based models highlights how factors like language- and/or domain-specific pre-training, few-shot sample size, and model architecture affect the detection and classification of potentially abusive clauses. Results show a large variability in performance for the different tasks and models, with the highest macro-F1 scores for the detection task ranging from 79% to 89% and micro-F1 scores up to 96%, while macro-F1 scores for the classification task range from 60% to 70% and micro-F1 scores from 64% to 80%. Notably, this is the first Spanish-language multi-label classification dataset for legal clauses, applying Chilean law and offering a comprehensive evaluation of Spanish-language models in the legal domain. Our work lays the ground for future research in method development for rarely considered legal analysis and potentially leads to practical applications to support consumers in Chile and Latin America as a whole.
@@
Consumer protection law, Abusive clauses, Natural language processing, Machine learning, Neural networks
@@ 1 Introduction
In the digital age, where a growing number of products and services are offered through online platforms, Terms of Service (ToS) have become increasingly relevant, since they usually constitute the step where the provider and the user become legally bound. However, ToS are long, difficult to read, sometimes not even available in the user’s mother tongue, and at the end of the day, pointless to review. Even if one could and would want to spend time reading them to use the service, there is no alternative but to accept them. It is impossible to change or negotiate a disadvantageous or even an abusive clause [19]. These problems of rational apathy and information asymmetry are present in most modern consumer contracts, but now they are amplified by the number and speed they occur through digital platforms. This constitutes a challenge for the regulator, the surveillance agencies, and the consumers [40].
In this context, the use of Artificial Intelligence in the legal field could be of tremendous impact, because it could automate the revision of these ToS and other consumer contracts, alerting the user and national services of the existence of potentially abusive clauses [36]. Recent developments have already led to promising systems, however, with few exceptions [60], these have mainly focused on the European Union [40, 8, 21, 53, 6, 23, 28]. Methods developed for other legal systems and languages, such as South American legislation developed in Spanish, are also necessary and should be included in the global research community’s attention.
Hence, our research objectives focus on developing Machine Learning (ML) for the automated identification of potentially abusive clauses in Chilean online ToS. Hereby, we build on work conducted on European online ToS [40, 53, 13, 23, 28] and update the methodology to recent developments in Natural Language Processing (NLP) [48, 22, 64, 41].
@@
  • We present a classification system categorizing 20 potentially abusive clauses into four groups of which we annotate three for machine learning tasks: 6 illegal, 6 dark, and 8 gray clauses.
  • We introduce the Chilean ToS dataset, comprising 50 legal documents in Spanish with a total of 5,209 annotated clauses. From these, we generate six datasets for detection and classification tasks. These datasets are imbalanced and can be used for binary or multi-class/-label classification.
@@
The rest of the paper is structured as follows. The next section formulates the problem from the Chilean legal perspective. Section 3 presents the proposed annotation system and introduces the Chilean Terms of Services corpus and its datasets. Section 4 explains ML methods and Section 5 describes the experimental setup and reports the results of our study. Section 6 discusses relevant prior research and provides an outlook, and Section 7 concludes.
@@ 2 Problem Formulation
In contracts of adhesion, that is, contracts, where one party sets the terms and conditions of the agreement and the counterpart is generally left only with the faculty to accept or reject the proposal as a whole [19], there is the risk that abusive clauses would be included. Legal systems refer to these types of clauses in various ways. In Chile, the Consumer Protection Law Act (Law Number 19,496, "Ley de Protección al Consumidor", recently updated and modified by the "Decreto con Fuerza de Ley Numero 3" from 2021, from now on LPC) does not define abusive clauses. Instead, it outlines the requirements that every consumer contract must fulfill and lists clauses that will be deemed invalid if introduced and that could potentially nullify the entire contract (Article 16 of the LPC).
@@
The Chilean National Service for Consumer Protection ("Servicio Nacional del Consumidor", SERNAC) has defined abusive clauses as those imposed by the provider of the goods or services in violation of the principles of equality and good faith, that result in a significant imbalance of contractual duties, in detriment of the weakest part, the consumer [56]. Examples of these clauses are the ones that allow the unilateral termination of the contract at the mere discretion of the provider, that is, without expression of cause or the fault of the consumer, that include disclaimers of liability even when the suppliers caused the damage, or that cause a reversal of the burden of proof, among many other abusive clauses.
@@
The core of these types of clauses lies in the power imbalance between the parties involved in the contract. One way to restore that balance is to report the existence of abusive clauses and eventually pursue legal action against providers who use them. However, the first requirement for this process is that consumers are aware of the existence of abusive clauses, which is not a simple task.
According to a survey conducted by SERNAC, over half of consumers are unable to identify potentially abusive clauses in contracts, while a third are completely unaware of their existence. [^1] https://www.sernac.cl/portal/604/w3-article-64918.html, (March 1st, 2022) This situation is exacerbated even further in the case of ToS, which are designed to be accepted just by one click in a matter of seconds [40].
@@
In this context, the general public could use specialized ML methods as a preventive tool to check if a ToS or other contracts include potentially abusive clauses. The model could swiftly accomplish the task, that would be arduous and time-consuming for a human. This model could automatically review the entire extension of the terms and raise an alarm if detects any clause that appears to be abusive. The primary objective of this research is to develop such methods.
@@
It is crucial to recognize that specific clauses can be interpreted in various ways, depending on the systematic analysis of the contract and the factual context in which they are applied. This causes interpretation problems that affect even specialized lawyers, posing a substantial challenge for ML algorithms. Furthermore, the judicial authority determines whether a clause is abusive. They will impose the appropriate consequences, which may include the annulment of the clause or contract, and eventually, the compensation of the suffered damages. Until the Supreme Court has settled the discussion, a clause is only potentially abusive. In other words, neither an AI method, lawyer nor even SERNAC can definitively determine whether a clause is abusive. Only the Supreme Court possesses the authority to affirm or deny a clause’s nature with legal certainty. Thus, our proposed ML methods and their results should be understood as a reference to guide further actions but are not intended to replace professional and technical advice. However, they can still serve as a very useful preventive tool that can raise awareness among consumers about the existence of abusive clauses and may support the work of surveillance agencies and lawyers who deal with such clauses.
3 Potentially Abusive Clauses
@@
In this study, we analyzed 50 Spanish-language ToS from companies that provide services in Chile. [^2] Academia.edu, Airbnb, Amazon, App Copec, Apple, Badoo, Bluexpress, Booking, Box, Canva, Despegar, Dropbox, eBay, Edreams, Evernote, Facebook, Fitbit, Google, Headspace, Instagram, Linkedln, Mercadolibre, Microsoft, MyHeritage, Nespresso, Netflix, Nike, Nintendo, Paris, Pokemon GO, Pullman Bus, Rappi, Ripley, Rovio, Skype, Skyscanner, Snapchat, Spotify, Starbucks, Tenpo, Tinder, TripAdvisor, Uber, Vimeo, WhatsApp, Wild Foods, World of Warcraft, X (Twitter), Yahoo, YouTube In total, this corresponds to 5,209 clauses, of which $20.3\%$ were identified as potentially abusive under at least one category. That means they were clauses that presented more than one reason that deemed them abusive, so the level of abusiveness is actually underrepresented. That is a surprisingly high level of abusive clauses since one could expect that standardized contracts used by internationally renowned corporations should comply with the law. Again this result reaffirms the necessity of developing tools that alert consumers of the existence of potentially abusive clauses.
@@
From the about 20% of clauses that were deemed abusive by our legal expert annotators, some of them were quite common. For example, we found 156 clauses in the corpus that limit the provider’s liability (ltd). On the other hand, there are extremely rare cases, such as the 10 clauses in the corpus that permit the provider the unilateral termination of the contract (ter). Furthermore, several clauses may be potentially abusive in multiple ways. For instance, out of the 117 clauses that impose external costs on the consumer (des risk), 34 were also against good faith (bfe). In the following sections, we describe our proposed categorization system and its type of potentially abusive clauses, the annotated clauses and finally the corpus statistics.
@@ 3.1 Categorization System
We developed our own categorization system for reviewing the ToS based on three factors. First, we used the legal classification Chilean doctrine has made of Abusive Clauses [19, 37, 4, 44, 52, 56]. Second, we reviewed the classification used in previous related work [40, 53]. Finally, we took the optimal methodology to identify the clauses into consideration. This work defines four categories of potentially abusive clauses for Chilean law: clauses that infringe formal requirements, and illegal, dark, or gray clauses. Only the last three of these four are predicted in our experiments, as the next section will clarify.
@@ 3.2 Formal Requirements of Clauses
Some norms of the LPC, like Articles 17 and 17A, specify different formal requirements that contracts must fulfill regardless of their type or content. For example, the minimum size of the letter that should be used, the language in which they should be written, and the obligation of providing a print copy of the contract, among others. Under these requirements, many of the ToS were at fault for very basic elements, for example, some of them were not available in Spanish [^3] Those were excluded from this study that focus on this language, or were not printable.
Even though relevant to consumers, these types of infringements of the law are not of special interest in our current research, since their identification is relatively straightforward and no specialized model is required for it. Instead, more traditional text parsers could be used to analyze, for example, the language of the ToS, its HTML formatting and font size, or the blocking of any print functionality via JavaScript. Therefore, they were excluded from the experimental analysis of this work, which focuses on the development and review of the performance of ML methods.
@@ 3.3 Illegal Clauses
Illegal Clauses refer to provisions that go against the explicit order of a norm to the detriment of the consumer. Therefore, even though they are included in a contract or ToS, they are not effective and should be considered as "not written", according to Articles 1,467 and 1,682 of the Civil Code ("Código Civil", CC) regarding the effect of annulment of clauses.
This category is not used by the Chilean doctrine or previous related work but is of our creation. We decided to separate this type of clause due to two reasons. First, because they constitute a relevant case of potentially abusive clauses (about 45% of them), and, secondly, because their characteristics make them both easy to recognize and simultaneously uniquely challenging for an AI system. They are easily recognizable because they are contrasted directly with a written norm that forbids them. However, the condition for their detection is the awareness of the legal rule being infringed, which constitutes the challenge of the category. The difficulty arises from the broad scope of potentially applicable norms because it is challenging to cover all the relevant ones since a contractual provision could potentially violate any legal rule of the Chilean system. Even though the ML model may successfully predict the lawfulness of clauses in an annotated dataset, the model is ultimately limited by its training scope. Therefore, it can be challenging to identify and group the applicable norms. For example, during the annotation of ToS, we found that a specific law was infringed by a ToS, that is, Law Number 20,720 regarding the bankruptcy regime and insolvency procedures. This was possible due to the annotators’ expert knowledge of bankruptcy law, but could not have been generally expected from a lawyer specialized in other fields. Similarly, maybe other clauses were unlawful but were not detected by the annotator for being outside of their scope of knowledge, and thus the ML model was not trained with the required data.
In the process, we noticed that some illegalities were more common than others, so we split them into subcategories. For example, the clauses that contravene provisions of the contractual liability regime of the CC included infringements in the rules for the determination of damages or the binding nature of contracts. Other usual illegal clauses were the ones that altered the jurisdiction of courts or prevented the consumer from acceding to one. Instead, they forced them to use informal conflict mechanism resolutions, violating procedural rules of the LPC, the Civil Procedure Code ("Código de Procedimiento Civil", CPC), and the Organic Code of Courts ("Código Orgánico de Tribunales, COT). Since many ToS come from international corporations, clauses referencing foreign law, that are not applicable in Chile (and sometimes even in direct conflict with national provisions), were also identified as illegal, to ensure local consumer rights are upheld.
@@ 3.4 Dark Clauses
Dark Clauses are the ones that, even though they are not directly illegal, are manifestly abusive. They are caused by the great power imbalance the provider holds, since the consumer can only agree with them or not sign the contract, i.e., not use the service. The provisions included in letters a), b), c) d), e) and h) of Article 16 LPC refer to this type of clauses [4]. These are, for example, the unilateral termination of the contract at the mere discretion of the provider; the arbitrary modification of the elements of the contract, including a unilateral price increase; preemptive limitation of liability clauses; and alteration of the burden of proof.
A commonality among these clauses is how evidently imbalanced they are, and for that reason, the Chilean legislator presumes their abusiveness. This means that, even though their identification as abusive still needs to be done by a court, the burden of proof lies with the supplier and not on the consumer. In other words, the provider that incorporated a dark clause into their contract needs to prove that, despite their appearance, they are not actually abusive. This constitutes an important advantage at the moment of presenting a claim, but a requirement for this step is still that consumers are aware of their existence and abusiveness. According to the studies conducted by SERNAC (see footnote 1), this is not a simple task for most people. Perhaps it is because of this that, despite the legislator identifying them as abusive, they continue to be widely used by providers.
Dark Clauses present three special characteristics that might be beneficial for the use of ML. First, the above-mentioned letters of Article 16 LPC constitute a well-defined list of clauses, that stands in contrast with the countless sources of unlawfulness that can make a clause illegal. Closed-set classification problems, such as this, are generally considered easier to solve than open-set classification like the Illegal Clauses problem. For instance, there is no need for open-set recognition [61]. However, the challenge of the variability of examples for one type of potentially abusive clause remains, i.e., they can be reformulated in different ways that are not recognizable anymore, for instance, using synonyms or a different grammatical structure. Second, since they are presumed abusive, there are fewer errors of interpretation and they are relatively clear. In contrast, Gray Clauses are more ambiguous, see Sec. 3.5. While the final decision is with the courts, their identification via ML would help raise awareness of the clauses’ existence among consumers. Third, since many Dark Clauses appear relatively frequently in the ToS that we reviewed, and thus, their patterns and specific formulations repeat as well, ML models may learn to predict these more easily. Conversely, the clauses that are less represented in the data may be harder to predict accurately.